[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Oct 30, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon Oct 30 14:06:56 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. GFE/NCP issues: - Transmit procedure in GFE Formatter Launcher is  
stalling again. This is a recurring problem, more random this time. TSB  
AWIPS Sysadmin Chris Mello will investigate.

2. Multiple problems with the AWIPS collaboration server: 1) "Unable to log  
into XMPP server" errors, 2) incorrect user role/name, 3) difficulties  
getting into collaboration room. NCF TT# 152191, opened in September,  
remains open but with MEDIUM priority, will be escalated to HIGH priority.  
TSB AWIPS Sysadmin Chris Mello will follow up.

3. Multiple problems with the NWPS wave modeling sytetm (hung processes,  
failure to untar bathymetry files, delayed/partial wave period and swell  
information, etc.) probably caused by reactivating post-processing and web  
site plotting, which is overloading the comp02 server. TSB is investigating  
whether TAFB has access to plots from the RWPS and whether WFOs are  
actually using the boundary conditions generated by NWPS runs in  
preparation for deactivating the post-processing.
HSU:
1. N-AWIPS issues: NOAA C-MAN station PLSF1/Pulaski Shoals, Florida is not  
plotting in NMAP.
http://www.ndbc.noaa.gov/station_page.php?station=plsf1
Location: 24.693 N, 82.773 W (west of Florida Keys)
TSB N-AWIPS Guru Matt Sardi will investigate
Other items:
1. AC Cooling Unit #2 in the Server room was alarming when we arrived  
[Monday] morning. There did not seem to be an issue with Unit #2 and the  
alarm was silenced. However, Unit #3 was not beeping but did show 2 alarms,  
both for High Head Pressure. This unit was also warmer than the other 2 at  
73 degrees vs. 68-69 degrees on the other two units. NHC Administrator  
Officer Jann Perez and William Martinez will follow up.
NHC RFCs:
1. RFC 255 - Implement a new script (nhc_ssuPsurge.pl) which allows the SSU  
(Storm Surge Unit) to submit the water levels file, submit a parallel  
Psurge run (instead of bothering HSU to do it), and submitting an internal  
Psurge run on WCOSS. To be implemented on Wednesday, November 1 at 1600Z  
(noon).

2. RFC 3852 - Install hardware upgrades for the HSU Atlantic desk and FEMA  
HLT N-AWIPS workstations. To be implemented on Monday, October 30 at 1300Z  
(9 am).

3. RFC 3855 - Add 5 additional TCP ports so GRB VPN users may access the  
WxConnect servers at NHC. Also, remove two VPN pools from the network ACLs  
(access control lists) allowing traffic to the GRB servers. To be  
implemented on Monday, October 30 at 1700Z (1 pm).

4. RFC 3858 - Perform final clean up and removal of NIS (network  
information service) authentication and server hardware at NHC because the  
NIS server at NHC is a single point of failure running on out-of-warranty  
hardware. Implementation to begin on Tuesday, October 31 at 1300Z (9 am)  
and should be completed by Friday, November 3.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171030/a6d0090d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list